# Towards Unsupervised Training of Matching-based Graph Edit Distance Solver via Preference-aware GAN

### v1

@@ Abstract
Graph Edit Distance (GED) is a fundamental graph similarity metric widely used in various applications. However, computing GED is an NP-hard problem. Recent state-of-the-art hybrid GED solver has shown promising performance by formulating GED as a bipartite graph matching problem, then leveraging a generative diffusion model to predict node matching between two graphs, from which both the GED and its corresponding edit path can be extracted using a traditional algorithm. However, such methods typically rely heavily on ground-truth supervision, where the ground-truth labels are often costly to obtain in real-world scenarios. In this paper, we propose GEDRanker, a novel unsupervised GAN-based framework for GED computation. Specifically, GEDRanker consists of a matching-based GED solver and introduces an interpretable preference-aware discriminator with an effective training strategy to guide the matching-based GED solver toward generating high-quality node matching without the need for ground-truth labels. Extensive experiments on benchmark datasets demonstrate that our GEDRanker enables the matching-based GED solver to achieve near-optimal solution quality without any ground-truth supervision.
@@ 1 Introduction
Graph Edit Distance (GED) is a widely used graph similarity metric [1, 2, 3] that determines the minimum number of edit operations required to transform one graph into another. It has broad applications in domains such as pattern recognition [4, 5] and computer vision [6, 7]. Figure 1(a) illustrates an example of an optimal edit path for transforming $G_{1}$ to $G_{2}$ with GED$(G_{1},G_{2})=4$. GED is particularly appealing due to its interpretability and its ability to capture both attribute-level and structural-level similarities between graphs. Traditional exact approaches for computing GED rely on A* search [8, 9, 10]. However, due to the NP-hardness of GED computation, these methods fail to scale to graphs with more than 16 nodes [9], making them impractical for large graphs.
To address the limitations of traditional algorithms, hybrid approaches that combine deep learning with traditional algorithms have been widely explored. Recent hybrid approaches have focused on reformulating GED as a bipartite graph matching problem. Specifically, GEDGNN [11] and GEDIOT [12] propose to train a node matching model that predicts a single node matching probability matrix, where top-$k$ edit paths can be extracted sequentially from the predicted matrices using a traditional algorithm. Later on, the current state-of-the-art GED solver, DiffGED [13], employs a generative diffusion model to predict top-$k$ high quality node matching probability matrices in parallel, which parallelized the extraction of top-$k$ edit paths. Notably, DiffGED has achieved near-optimal solution quality, closely approximating the ground-truth GED with a remarkable computational efficiency. Unfortunately, all existing hybrid GED solvers heavily rely on supervised learning, requiring ground-truth node matching matrices and GED values, which are often infeasible to obtain in practice due to the NP-hard nature of GED computation.
This limits the usability of existing GED methods in real-world scenarios. Therefore, developing an unsupervised training framework for GED solvers is crucial for enhancing real-world applicability. Despite its importance, this direction remains largely underexplored.
@@
Contributions. To the best of our knowledge, GEDRanker is the first unsupervised learning framework for GED computation, providing a scalable and label-free alternative to supervised methods. Experimental results on benchmark datasets show that the current state-of-the-art supervised matching-based GED solver, when trained under our unsupervised framework, achieves near-optimal solution quality comparable to supervised training with the same number of epochs, and outperforms all other supervised learning methods and traditional GED solvers.
Figure 1: (a) An optimal edit path for converting $G_{1}$ to $G_{2}$ with GED$(G_{1},G_{2})=4$. (b) An optimal node matching matrix from which an optimal edit path can be derived.
@@ 3.1 Problem Formulation
(2) Set all values in the $v$-th row and $u$-th column of $\hat{\pi}$ to $\infty$- ∞.
@@ 3.2 Supervised Diffusion-based Node Matching Model
Figure 2: An overview of GEDRanker. For each training step, given a pair of training graphs, we maintain a record of the current best node matching matrix $\bar{\pi}$ and the node matching matrix obtained from the previous training step $\pi_{last}$. A noisy matching $\pi^{t}$ is sampled at a random diffusion time step $t$ and denoised by the denoising network $g_{\phi}$ to produce node matching scores. The resulting matching probability matrix $\hat{\pi}_{g_{\phi}}$ is obtained via Gumbel-Sinkhorn and greedily decoded to $\pi_{g_{\phi}}$. The preference-aware discriminator $D_{\theta}$ is trained to rank the preference over $\bar{\pi}$, $\pi_{\text{last}}$, and $\hat{\pi}_{g_{\phi}}$. Next, $g_{\phi}$ is trained to recover $\bar{\pi}$ and maximize the preference score $D_{\theta}(G_{1},G_{2},\hat{\pi}_{g_{\phi}})$ Finally, the record is updated by $\pi_{g_{\phi}}$.
@@ 4.1 Unsupervised Training Strategy
In practice, the ground-truth $\pi^{*}$ is often unavailable, making it impractical to directly apply the supervised training strategy. A modified training strategy is to store a randomly initialized binary constrained node matching matrix as the current best solution $\bar{\pi}$ for each training graph pair. At each training step, for a given pair of graphs, we sample a random $t$, and sample a noisy matching matrix $\pi^{t}$ from $q(\pi^{t}|\pi^{0})$, with $\pi^{0}=\bar{\pi}$, then $g_{\phi}$ is trained to recover $\bar{\pi}$ from $\pi^{t}$ by minimizing $\mathcal{L}_{rec(\bar{\pi})}$. If the node matching matrix $\pi_{g_{\phi}}$ decoded from $\hat{\pi}_{g_{\phi}}=\sigma(g_{\phi}(G_{1},G_{2},\pi^{t},t))$ yields a shorter edit path (i.e., $c(G_{1},G_{2},\pi_{g_{\phi}})<c(G_{1},G_{2},\bar{\pi})$), then $\bar{\pi}$ is updated to $\pi_{g_{\phi}}$. This guarantees that $g_{\phi}$ progressively finds a better solution during training, while recovering the current best solution.
@@
$\nabla_{\phi}J(\phi)=\mathbb{E}_{\pi_{g_{\phi}}}[(c(G_{1},G_{2},\pi_{g_{\phi}})-b)\nabla_{\phi}\log{\sigma(g_{\phi}(G_{1},G_{2},\pi^{t},t))}]$ (4)
where $b$ is a baseline used to reduce gradient variance. This method adjusts the parameters of $g_{\phi}$ to increase or decrease the matching probability of each node pair based on the resulting edit path length, encouraging the model to generate node matching probability matrices that are more likely to yield shorter edit paths.
@@ 4.2 Discriminator
Preference-aware Training Objective. To ensure that $D_{\theta}$ learns the correct preference, we introduce a preference-aware discriminator. A node matching matrix with a shorter edit path should be preferred ()succeeds($\succ$)( ≻ ) over one with a longer edit path. The objective of the preference-aware discriminator is to guarantee that the score assigned to the preferred node matching matrix is ranked higher than that of the less preferred one. Specifically, at each training step, given a pair of graphs, we have the current best node matching matrix $\bar{\pi}$ and the predicted node matching probability matrix $\hat{\pi}_{g_{\phi}}=S(g_{\phi}(G_{1},G_{2},\pi^{t},t))$. The preference-aware discriminator is trained to minimize the following Bayes Personalized Ranking (BPR) loss [28]:
@@
Discriminator Architecture. To effectively capture the dependencies among node pairs and distinguish the influence of each predicted node matching on the resulting edit path, our $D_{\theta}$ leverages GIN [29] and Anisotropic Graph Neural Network (AGNN) [30, 31, 32] to compute the embeddings of each node pair, then estimates the preference score directly based on these embeddings. Due to the sapce limitation, more details about the architecture of $D_{\theta}$ can be found in Appendix A.3 and C.3.
@@ 5.1 Experimental Settings
Datasets. We conduct experiments on three widely used real world datasets: AIDS700 [17], Linux [33, 17], and IMDB [17, 34]. Each dataset is split into $60\%$, $20\%$ and $20\%$ as training graphs, validation graphs and testing graphs, respectively. We construct training, validation, and testing graph pairs, and generate their corresponding ground-truth GEDs and node matching matrices following the strategy described in [11]. More details of the datasets can be found in Appendix B.1.
@@
Evaluation Metrics. We evaluate the performance of each model using the following metrics: (1) Mean Absolute Error (MAE) measures average absolute error between the predicted GED and the ground-truth GED; (2) Accuracy measures the proportion of test pairs whose predicted GED equals the ground-truth GED; (3) Spearman’s Rank Correlation Coefficient ($\rho$), and (4) Kendall’s Rank Correlation Coefficient ($\tau$) both measure the matching ratio between the ranking results of the predicted GED and the ground-truth GED for each query test graph; (5) Precision at $10/20$ ($p$@$10$, $p$@$20$) measures the proportion of predicted top-$10/20$ most similar graphs that appear in the ground-truth top-$10/20$ similar graphs for each query test graph. (6) Time (s) measures the average running time of each test graph pair.
5.2 Main Results
@@
Datasets Models Type MAE $\downarrow$ Accuracy $\uparrow$ $\rho$uparrow $\tau$uparrow $p$@$10$uparrow $p$@$20$uparrow Time(s) $\downarrow$
@@
GENN-A* SL $0.632$ $61.5\%$ $0.903$ $0.815$ $85.6\%$ $88\%$ $2.98919$
MATA* SL $0.838$ $58.7\%$ $0.8$ $0.718$ $73.6\%$ $77.6\%$ 0.00487
GEDGNN SL $1.098$ $52.5\%$ $0.845$ $0.752$ $89.1\%$ $88.3\%$ $0.39448$
@@ 5.3 Ablation Study
Datasets Models MAE $\downarrow$ Accuracy $\uparrow$ $\rho$uparrow $\tau$uparrow $p$@$10$uparrow $p$@$20$uparrow
@@
GEDRanker (RL) $0.04$ $98\%$ $0.99$ $0.984$ $99.2\%$ $99.2\%$
GEDRanker (GED) $0.017$ $99.2\%$ $0.995$ $0.992$ $99.4\%$ $99.6\%$
GEDRanker (Hinge) $0.002$ $99.9\%$ $0.999$ $0.999$ $100\%$ $100\%$
@@
Ranking Loss. For the preference-aware discriminator, we adopt the BPR ranking loss. To evaluate the effect of alternative ranking loss, we introduce a variant model, GEDRanker (Hinge), which replaces BPR with a Hinge loss to rank the preference over different node matching matrices $\pi$. Following the same setting as in Equation 8, the Hinge loss is computed with a margin of $1$ when the two node matching matrices yield different edit path lengths, and with a margin of 00 when they result in the same edit path length. In the latter case, the loss is computed in both directions to encourage their predicted scores to be as similar as possible. The detailed Hinge loss is shown in Appendix A.4.

### v2

@@ Abstract
Graph Edit Distance (GED) is a fundamental graph similarity metric widely used in various applications. However, computing GED is an NP-hard problem. Recent state-of-the-art hybrid GED solver has shown promising performance by formulating GED as a bipartite graph matching problem, then leveraging a generative diffusion model to predict node matching between two graphs, from which both the GED and its corresponding edit path can be extracted using a traditional algorithm. However, such methods typically rely heavily on ground-truth supervision, where the ground-truth node matchings are often costly to obtain in real-world scenarios. In this paper, we propose GEDRanker, a novel unsupervised GAN-based framework for GED computation. Specifically, GEDRanker consists of a matching-based GED solver and introduces an interpretable preference-aware discriminator. By leveraging preference signals over different node matchings derived from edit path lengths, the discriminator can guide the matching-based solver toward generating high-quality node matching without the need for ground-truth supervision. Extensive experiments on benchmark datasets demonstrate that our GEDRanker enables the matching-based GED solver to achieve near-optimal solution quality without any ground-truth supervision. The source code is available at https://github.com/piupiupiuu/GEDRanker.
@@ 1 Introduction
Graph Edit Distance (GED) is a widely used graph similarity metric [1, 2, 3] that determines the minimum number of edit operations required to transform one graph into another. It has broad applications in domains such as pattern recognition [4, 5] and computer vision [6, 7]. Figure 1(a) illustrates an example of an optimal edit path for transforming $G_{1}$ to $G_{2}$ with $(G_{1},G_{2})=4. GED$ is particularly appealing due to its interpretability and its ability to capture both attribute-level and structural-level similarities between graphs. Traditional exact approaches for computing GED rely on A* search [8, 9, 10]. However, due to the NP-hardness of GED computation, these methods fail to scale to graphs with more than 16 nodes [9], making them impractical for large graphs.
To address the limitations of traditional algorithms, hybrid approaches that combine deep learning with traditional algorithms have been widely explored. Recent hybrid approaches have focused on reformulating GED as a bipartite graph matching problem. Specifically, GEDGNN [11] and GEDIOT [12] propose to train a node matching model that predicts a single node matching probability matrix, where top-$k$ edit paths can be extracted sequentially from the predicted matrices using a traditional algorithm. Later on, the current state-of-the-art GED solver, DiffGED [13], employs a generative diffusion model to predict top-$k$ high quality node matching probability matrices in parallel, which parallelized the extraction of top-$k$ edit paths. Notably, DiffGED has achieved near-optimal solution quality, closely approximating the ground-truth GED with a remarkable computational efficiency. Unfortunately, all existing hybrid GED solvers heavily rely on supervised learning, requiring ground-truth node matching matrices and GED values, which are often infeasible to obtain in practice due to the NP-hard nature of GED computation. This limits the usability of existing GED methods in real-world scenarios. Therefore, developing an unsupervised training framework for GED solvers is crucial for enhancing real-world applicability. Despite its importance, this direction remains largely underexplored.
@@
Contributions. To the best of our knowledge, GEDRanker is the first framework that enables unsupervised training of supervised neural methods for GED computation, offering a scalable and label-free alternative to supervised approaches. Experimental results on benchmark datasets show that the current state-of-the-art supervised matching-based GED solver, when trained under our unsupervised framework, achieves near-optimal solution quality comparable to supervised training with the same number of epochs, and outperforms all other supervised learning methods and traditional GED solvers.
[…]
@@ 3.1 Problem Formulation
(2) Set all values in the $v$-th row and $u$-th column of $\hat{\pi}$ to -$\infty$.
@@ 3.2 Supervised Diffusion-based Node Matching Model
Figure 2: An overview of GEDRanker. For each training step, given a pair of training graphs, we maintain a record of the current best node matching matrix $\bar{\pi}$ and the node matching matrix obtained from the previous training step $\pi_{last}$. A noisy matching $\pi^{t}$ is sampled at a random diffusion time step $t$ and denoised by the denoising network $g_{\phi}$ to produce node matching scores. The resulting matching probability matrix $\hat{\pi}_{g_{\phi}}$ is obtained via Gumbel-Sinkhorn and greedily decoded to $\pi_{g_{\phi}}$. The preference-aware discriminator $D_{\theta}$ is trained to learn a preference ordering over $\bar{\pi}$, $\pi_{\text{last}}$, and $\hat{\pi}_{g_{\phi}}$. Next, $g_{\phi}$ is trained to recover $\bar{\pi}$ and maximize the preference score $D_{\theta}(G_{1},G_{2},\hat{\pi}_{g_{\phi}})$ . Finally, the record is updated by $\pi_{g_{\phi}}$.
@@ 4.1 Unsupervised Training Strategy
In practice, the ground-truth $\pi^{*}$ is often unavailable, making it impractical to directly apply the supervised training strategy. Assuming the same set of training graphs as in the supervised learning setting, a modified training strategy is to store a randomly initialized binary constrained node matching matrix as the current best solution $\bar{\pi}$ for each training graph pair. At each training step, for a given pair of graphs, we sample a random $t$, and sample a noisy matching matrix $\pi^{t}$ from $q(\pi^{t}|\pi^{0})$, with $\pi^{0}=\bar{\pi}$, then $g_{\phi}$ is trained to recover $\bar{\pi}$ from $\pi^{t}$ by minimizing $\mathcal{L}_{rec(\bar{\pi})}$. If the node matching matrix $\pi_{g_{\phi}}$ decoded from $\hat{\pi}_{g_{\phi}}=\sigma(g_{\phi}(G_{1},G_{2},\pi^{t},t))$ yields a shorter edit path (i.e., $c(G_{1},G_{2},\pi_{g_{\phi}})<c(G_{1},G_{2},\bar{\pi})$), then $\bar{\pi}$ is updated to $\pi_{g_{\phi}}$. This guarantees that $g_{\phi}$ progressively finds a better solution during training, while recovering the current best solution.
@@
$\nabla_{\phi}J(\phi)=\mathbb{E}_{\pi_{g_{\phi}}}[(c(G_{1},G_{2},\pi_{g_{\phi}})-b)\nabla_{\phi}\log{\sigma(g_{\phi}(\pi_{g_{\phi}}|G_{1},G_{2},\pi^{t},t))}]$ (4)
where $c(G_{1},G_{2},\pi_{g_{\phi}})$ is the cost of the edit path computed according to the Edit Path Generation procedure in Section 3.1, and $b$ is a baseline used to reduce gradient variance. This method adjusts the parameters of $g_{\phi}$ to increase or decrease the matching probability of each node pair based on the resulting edit path length, encouraging the model to generate node matching probability matrices that are more likely to yield shorter edit paths.
@@ 4.2 Discriminator
Preference-aware Training Objective. To ensure that $D_{\theta}$ learns the correct preference, we introduce a preference-aware discriminator. A node matching matrix with a shorter edit path should be preferred $(\succ)$ over one with a longer edit path. The objective of the preference-aware discriminator is to guarantee that the score assigned to the preferred node matching matrix is ranked higher than that of the less preferred one. Specifically, at each training step, given a pair of graphs, we have the current best node matching matrix $\bar{\pi}$ and the predicted node matching probability matrix $\hat{\pi}_{g_{\phi}}=S(g_{\phi}(G_{1},G_{2},\pi^{t},t))$. The preference-aware discriminator is trained to minimize the following Bayes Personalized Ranking (BPR) loss [28]:
@@
Discriminator Architecture. To effectively capture the dependencies among node pairs and distinguish the influence of each predicted node matching on the resulting edit path, our $D_{\theta}$ leverages GIN [29] and Anisotropic Graph Neural Network (AGNN) [30, 31, 32] to compute the embeddings of each node pair, then estimates the preference score directly based on these embeddings. Due to the space limitation, more details about the architecture of $D_{\theta}$ can be found in Appendix A.3 and C.3.
@@ 5.1 Experimental Settings
Datasets. We conduct experiments on three widely used real world datasets: AIDS700 [17], Linux [33, 17], and IMDB [17, 34]. Each dataset is split into $60\%$, $20\%$ and $20\%$ as training graphs, validation graphs and testing graphs, respectively. We construct training, validation, and testing graph pairs, and generate their corresponding ground-truth GEDs and node matching matrices for evaluation following the strategy described in [11]. More details of the datasets can be found in Appendix B.1.
@@
Evaluation Metrics. We evaluate the performance of each model using the following metrics: (1) Mean Absolute Error (MAE) measures average absolute error between the predicted GED and the ground-truth GED; (2) Accuracy measures the proportion of test pairs whose predicted GED equals the ground-truth GED; (3) Spearman’s Rank Correlation Coefficient ($\rho$), and (4) Kendall’s Rank Correlation Coefficient ($\tau$) both measure the matching ratio between the ranking results of the predicted GED and the ground-truth GED for each query test graph; (5) Precision at $10/20$ (pp@$10$, pp@2020) measures the proportion of predicted top-$10/20$ most similar graphs that appear in the ground-truth top-$10/20$ similar graphs for each query test graph. (6) Time (s) measures the average running time of each test graph pair.
@@
Datasets Models Type MAE $\downarrow$ Accuracy $\uparrow$ $\rho$ $\uparrow$ $\tau$ $\uparrow$ pp@$10$ $\uparrow$ pp@2020 $\uparrow$ Time(s) $\downarrow$
@@
GENN-A* SL $0.632$ $61.5\%$ $0.903$ $0.815$ $5.6\% 88$$8\%$ $2.98919$
MATA* SL $0.838$ $58.7\%$ $0.8$ $0.718$ $3.6\% 77$$7.6\%$ 0.00487
GEDGNN SL $1.098$ $52.5\%$ $0.845$ $0.752$ $9.1\% 88$$8.3\%$ $0.39448$
@@ 5.3 Ablation Study
Datasets Models MAE $\downarrow$ Accuracy $\uparrow$ $\rho$ $\uparrow$ $\tau$ $\uparrow$ pp@$10$ $\uparrow$ pp@2020 $\uparrow$
@@
GEDRanker (RL) $0.04$ $98\%$ $0.99$ $0.984$ $\% 99.299.2\%$
GEDRanker (GED) $0.017$ $99.2\%$ $0.995$ $0.992$ $9.4\% 99$$9.6\%$
GEDRanker (Hinge) $0.002$ $99.9\%$ $0.999$ $0.999$ $\% 100100\%$
@@
Ranking Loss. For the preference-aware discriminator, we adopt the BPR ranking loss. To evaluate the effect of alternative ranking loss, we introduce a variant model, GEDRanker (Hinge), which replaces BPR with a Hinge loss to rank the preference over different node matching matrices $\pi$. Following the same setting as in Equation 8, the Hinge loss is computed with a margin of $1$ when the two node matching matrices yield different edit path lengths, and with a margin of 0 when they result in the same edit path length. In the latter case, the loss is computed in both directions to encourage their predicted scores to be as similar as possible. The detailed Hinge loss is shown in Appendix A.4.